# The Collection of a Human Robot Collaboration Dataset for Cooperative Assembly in Glovebox Environments

### v1

@@ 4.2 Results
Experiment A. As indicated in Table 2, the pretraining experiments posed significant challenges across all models and datasets. We identify two primary causes for these difficulties. First, the emphasis on real-time models led to the selection of lower-capacity models, which are potentially less suited for pretraining tasks. Second, the HaDR dataset (18), being the largest, may have been the only dataset with sufficient diversity and scale to facilitate effective pretraining. While these observations regarding the necessity of model and dataset scale align with current best practices, they conflict with the requirements of real-time segmentation tasks.
Table 2: Transferability Assessment of Prior Work: Models are pretrained to achieve 0.8 IOU on their respective pretraining validation sets. Subsequently, these models are fine-tuned for 20 epochs using a random 10% sample from the HAGS dataset. A ‘-’ indicates a near-zero score.
ID OOD - Hands + Replaced GS
Model IOU $\uparrow$ ECE $\downarrow$ IOU $\uparrow$ ECE $\downarrow$
UNet+ensemble+dropout (HaDR) 0.4475 0.0039 0.2984 0.0089
UNet+ensemble+dropout (HRC) - - - -
UNet+ensemble+dropout (WH) - - - -
BiSeNetv2+ensemble+dropout (HaDR) 0.0025 0.0078 0.0015 0.0038
BiSeNetv2+ensemble+dropout (HRC) 0.1142 0.0155 0.0737 0.0167
BiSeNetv2+ensemble+dropout (WH) - - - -
@@
Table 3: Fine-tuning UNet and BiSeNetv2 from scratch on the HAGS ID training set and evaluating OOD performance on the ungloved test split (OOD-Hands). MobileSAM is fine-tuned from its published checkpoint. Reported metrics include Intersection over Union (IOU), Expected Calibration Error (ECE), and Predictive Entropy (PE). The reported is the difference between the OOD PE and ID PE: $\Delta PE=PE_{OOD}-PE_{ID}$.
ID OOD - Hands
Model IOU $\uparrow$ ECE $\downarrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$
UNet 0.8003 0.0034 0.5663 $0$0.0050
UNet+dropout 0.7897 0.0036 0.5709 0.0095 -0.0041
UNet+ensemble 0.8017 0.0043 0.5936 $0$0.0032
UNet+ensemble+dropout 0.8019 0.0043 0.5753 $0$0.0027
BiSeNetv2 0.7304 0.0052 0.5410 0.0103 -0.0040
BiSeNetv2+dropout 0.7320 0.0059 0.5513 0.0102 -0.0001
BiSeNetv2+ensemble 0.7427 0.0056 0.5321 0.0120 -0.0038
BiSeNetv2+ensemble+dropout 0.7384 0.0058 0.4991 0.0128 -0.0037
MobileSAM 0.5622 0.4020 0.5126 0.4059 0.0000
MobileSAM+dropout 0.5486 0.4011 0.5220 0.4057 0.0002
MobileSAM+ensemble 0.5168 0.4012 0.5128 0.4210 0.0009
MobileSAM+ensemble+dropout 0.4974 0.4004 0.5315 0.4223 0.0006
@@
Experiment B. In Table 3, we present the performance metrics on the ID test set, along with an OOD split focused exclusively on hands. The results for additional OOD splits, as elaborated in Section 4, are detailed in Table 4. Our analysis reveals that the primary source of variability in IoU scores arises from the choice of evaluation set rather than other factors such as the application of dropout or variations in model architecture. Notably, model ensembling leads to improved IoU scores, particularly evident in the replaced green screen dataset.
Table 4: An extension of Table 3 that reports the results on all of the OOD splits.
OOD - Hands OOD - Replaced GS OOD - Hands + Replaced GS
Model IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$
UNet 0.5663 $0$0.0050 0.5628 0.0114 -0.0041 0.5694 0.0104 -0.0046
UNet+dropout 0.5709 0.0095 -0.0041 0.5556 0.0115 -0.0028 0.5802 0.0098 -0.0037
UNet+ensemble 0.5936 $0$0.0032 0.6458 0.0095 -0.0022 0.6056 0.0099 -0.0024
UNet+ensemble+dropout 0.5753 $0$0.0027 0.6484 0.0092 -0.0015 0.6125 0.0096 -0.0028
BiSeNetv2 0.5410 0.0103 -0.0040 0.5447 0.0114 0.0016 0.5335 0.0111 -0.0021
BiSeNetv2+dropout 0.5513 0.0102 -0.0001 0.5495 0.0127 0.0029 0.5392 0.0116 0.0016
BiSeNetv2+ensemble 0.5321 0.0120 -0.0038 0.5857 0.0109 -0.0011 0.5409 0.0119 -0.0025
BiSeNetv2+ensemble+dropout 0.4991 0.0128 -0.0037 0.5622 0.0121 0.0017 0.5429 0.0123 -0.0027
MobileSAM 0.5126 0.4059 0.0000 0.5136 0.4067 0.0002 0.5196 0.4076 0.0004
MobileSAM+dropout 0.5220 0.4057 0.0002 0.5188 0.4065 0.0002 0.5263 0.4074 0.0004
MobileSAM+ensemble 0.5128 0.4210 0.0009 0.4802 0.4171 0.0013 0.4792 0.4070 0.0013
MobileSAM+ensemble+dropout 0.5315 0.4223 0.0006 0.5212 0.4193 0.0007 0.5112 0.4086 0.0012
@@ 5 Discussion, Conclusions, and Future Work
There are several areas in which this work can be improved. One significant enhancement would be to increase the diversity and quantity of images and labels, particularly by increasing the training set size to be comparable to other industrial domain datasets. Although the test subjects were diverse, the size of the studies limited the range of diversity that could be included. Future work should further improve diversity with a larger participant pool that includes various age ranges, genders, and skin tones. We currently employ two static camera angles, which, while providing some diversity, resulted in hands being predominantly located in predictable, centered portions of the image, as indicated by the pixel occupancy heat maps (see Figure 3). This is a noted, but not significant, limitation since the rigid form of the glovebox limits operators to working in a known and relatively small reachable workspace. The study does not utilize depth sensors to acquire accompanying depth data, nor does it record robot trajectory information. Although both of which could be valuable for advancing other human-robot interaction (HRI) applications, their exclusion simplifies the experimental set-up while mirroring glovebox configurations in the real world. A final possible limitation is that the participants utilized a single brand of glove. Although gloves are typically similar in color they can vary slightly; and more importantly, they tend to yellow with age.
@@
Acknowledgments and Disclosure of Funding
This manuscript has been approved for unlimited release and has been assigned LA-UR-24-25500. This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This work was supported by the Laboratory Directed Research and Development program of LANL under project number 20210043DR and C1582/CW8217. LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001).
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Please refer to Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Data is hosted in the public repository at https://dataverse.tdl.org/dataset.xhtml?persistentId=doi:10.18738/T8/85R7KQ and our code is available at https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset.
    2. (b)
[…]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] The IRB exemption document is available here: https://dataverse.tdl.org/file.xhtml?fileId=599918&version=1.0 and the instructions and risks are outlined here: https://dataverse.tdl.org/file.xhtml?fileId=599917&version=1.0
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] These instructions and warnings are available at https://dataverse.tdl.org/file.xhtml?fileId=599917&version=1.0
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] The hourly wage was \$0.00 per hour barring winning a raffled giftcard. These documents are available at https://dataverse.tdl.org/file.xhtml?fileId=599918&version=1.0 and https://dataverse.tdl.org/file.xhtml?fileId=599917&version=1.0
Appendix A Appendix
A.1 HAGS Datasheet
This section introduced our Hand and Glove Segmentation Dataset (HAGS) datasheet [17]. The HAGS dataset was collected by the Nuclear and Applied Robotics Group and funded by Los Alamos National Laboratory.
Hosting, Long Term Preservation, and Maintenance: The dataset will be maintained by Selma Wanna (Email: slwanna@utexas.edu) and hosted on the Texas Data Repository  [^7] https://dataverse.tdl.org/.
@@ A.1 HAGS Datasheet
The dataset consists of 191 videos, totaling approximately 9 hours of footage, along with 1728 annotated in-distribution and out-of-distribution frames, all originating from 12 diverse participants. Furthermore, the dataset emphasizes participants and procedures with various characteristics, such as differing skin tones, arm orientations, camera perspectives, and task execution methods. The two joint assembly tasks within this dataset include the assembly of a block tower and the disassembly of a tool box. In all, the diversity presented enables the dataset to cover a broad range of real-world scenarios.
@@
There are 16 videos per test subject, from which sampled frames and annotations were created over 2 experiments, 4 variables, 2 camera angles, and 12 subjects.
The dataset includes 12 diverse participants and is self-contained; it does not include external sources or assets.
@@
There are labels provided only for the images we have sampled. It is possible for other users to sample more images using the provided script  [^8] https://github.com/sanatnair/Glovebox_Segmentation_Dataset_Tools. Thereafter, software like LabelStudio [34] can assist in creating the new images’ ground truths.
@@
The dataset is self-contained barring the instances with green screens with overlaid images. These images were sourced from Google images using a script [^9] https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset/blob/main/data/convert_scripts/replace_green.py. There is no guarantee that the specific images will remain constant over time. There is not an archival version of this portion of the dataset that we can distribute.
@@
An ethical review process was conducted. An IRB exemption was granted for this study. This information is viewable online  [^10] https://dataverse.tdl.org/file.xhtml?fileId=599918&version=1.0.
@@
All participants received an informed consent form [^11] https://dataverse.tdl.org/file.xhtml?fileId=599917&version=1.0 and were aware of our study and data collection.
@@
The software for labeling the data is available from LabelStudio [34].
Frames are labeled as left and right hand with LabelStudio software [34] and the MobileSAM [40] backend. Models were trained on frames compressed to 256x256 px. Ground truth masks and original footage data is available, while raw unprocessed data is private.
@@
This repository  [^12] https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset/tree/main links to our public repository, which will contain a running list of associated papers or publications.
@@
A.2 Overview of Hand Segmentation Datasets
@@
GTEA [15] 2011 C GoPro Cooking Kitchen
ADL [27] 2012 C GoPro Daily Life Home
EDSH [22] 2013 C - Daily Life Home / Outdoors
Interactive Museum [2] 2014 C - Gesture Museum
EgoHands [1] 2015 C Google Glass Manipulation Social Setting
Maramotti [3] 2015 C - Gesture Museum
UNIGE Hands [4] 2015 C GoPro Hero3+ Daily Life Various Daily Locations
GUN-71 [29] 2015 CD Creative Senz3D Daily Life Home
RGBD Egocentric Action [35] 2015 CD Creative Senz3D Daily Life Home
Fingerwriting in mid-air [8] 2016 CD Creative Senz3D Writing Office
Ego-Finger [19] 2016 C - Gesture Various Daily Locations
ANS able-bodied [25] 2016 C Looxie 2 Daily Life Home
UT Grasp [6] 2016 C GoPro Hero2 Manipulation -
GestureAR [26] 2017 C Nexus 6 and Moto G3 Gesture Various Backgrounds
EgoGesture [37] 2017 C - Gesture -
Egocentric hand-action [38] 2017 D Softkinetic DS325 Gesture Office
BigHand2.2M [39] 2017 D Intel RealSense SR300 Gesture/Pose -
Desktop Action [7] 2018 C GoPro Hero 2 Daily Life Office
Epic Kitchens [11] 2018 C GoPro Cooking Kitchen
FPHA [16] 2018 CD Intel RealSense SR300 Gesture / Pose / Daily life Home
EYTH [21] 2018 C - - -
HandOverFace [21] 2018 C - - -
EGTEA+ [23] 2018 C SMI wearable eye-tracker Cooking Kitchen
THU-READ [33] 2018 CD Primesense Carmine Daily Life Home
EgoGesture [41] 2018 CD Intel RealSense SR300 Gesture -
HOI [20] 2018 D Kinect V2 Manipulation -
EgoDaily [10] 2019 C GoPro Hero5 Daily Life Various Daily Locations
ANS SCI [24] 2019 C GoPro Hero4 Daily Life Home
KBH [36] 2019 C HTC Vive Manipulation / Keyboard Office
WorkingHands [32] 2019 CD Kinect V2 Assembly Industrial
Freihand [42] 2019 C - Grasp / Pose Green Screen and Outdoors
MECCANO [28] 2020 C Intel RealSense SR300 Assembly Desk
HRC [30] 2022 C iPhone 11 Pro Assembly Industrial
HaDR [18] 2023 CD Realsense L515 Manipulation Industrial

### v2

@@ 4.2 Results
Experiment A. As shown in Figure 3, the transfer learning experiments presented a significant challenge in all models and datasets evaluated. These challenges are attributed to our focus on real-time applications, which required the use of lower-capacity models. The nature of our task deviates from others, e.g., offline image classification tasks, which may favor the prevailing paradigm of large-scale models and datasets.
Despite notable variance across runs, the HRC dataset emerged as a promising pretraining candidate, particularly in conjunction with the UNet architecture (30 million parameters), which demonstrated superior trainability compared to the BiSeNetv2 model (5 million parameters). However, this variability complicates definitive conclusions regarding the feasibility and necessity of pretraining for our downstream tasks. Specifically, for real-time hand and glove segmentation, the collection of the HRC dataset was essential. No alternative pretraining dataset was able to match the trained-from-scratch fine-tuning performance achieved by the real-time models, underscoring the critical role of task-specific dataset design. Despite the variability, there is some evidence suggesting a modest improvement in downstream task performance when 70% of the HAGS dataset is incorporated into the fine tuning process.
The HaDR dataset (Grushko et al., 2023), the largest pretraining dataset evaluated in this study, produced suboptimal results. This outcome suggests that, given the constraints of small model sizes, curating a smaller, more representative subset of the HaDR dataset could have been more effective than utilizing the dataset in its entirety. However, achieving such a subset requires significant engineering effort, involving extensive trial and error, to identify the most informative samples for the target task.
These findings highlight the nuanced trade-offs between dataset size, model capacity, and downstream performance in the context of real-time applications, where computational constraints play a critical role. Further exploration into dataset engineering and transfer learning stability is warranted to refine these insights.
[…]
@@
Table 2: Fine-tuning UNet and BiSeNetv2 from scratch on the HAGS ID training set and evaluating OOD performance on the ungloved test split (OOD-Hands). MobileSAM is fine-tuned from its published checkpoint. Reported metric include Intersection over Union (IOU), Expected Calibration Error (ECE), and Predictive Entropy (PE). The reported is the difference between the OOD PE and ID PE: $\Delta PE=PE_{OOD}-PE_{ID}$.
ID OOD - Hands OOD - Replaced GS OOD - Hands + Replaced GS
Model IOU $\uparrow$ ECE $\downarrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$ IOU $\uparrow$ ECE $\downarrow$ $\Delta$PE $\uparrow$
UNet 0.8003 0.0034 0.5663 $0$0.0050 0.5628 0.0114 -0.0041 0.5694 0.0104 -0.0046
UNet+dropout 0.7897 0.0036 0.5709 0.0095 -0.0041 0.5556 0.0115 -0.0028 0.5802 0.0098 -0.0037
UNet+ensemble 0.8017 0.0043 0.5936 $0$0.0032 0.6458 0.0095 -0.0022 0.6056 0.0099 -0.0024
UNet+ensemble+dropout 0.8019 0.0043 0.5753 $0$0.0027 0.6484 0.0092 -0.0015 0.6125 0.0096 -0.0028
BiSeNetv2 0.7304 0.0052 0.5410 0.0103 -0.0040 0.5447 0.0114 0.0016 0.5335 0.0111 -0.0021
BiSeNetv2+dropout 0.7320 0.0059 0.5513 0.0102 -0.0001 0.5495 0.0127 0.0029 0.5392 0.0116 0.0016
BiSeNetv2+ensemble 0.7427 0.0056 0.5321 0.0120 -0.0038 0.5857 0.0109 -0.0011 0.5409 0.0119 -0.0025
BiSeNetv2+ensemble+dropout 0.7384 0.0058 0.4991 0.0128 -0.0037 0.5622 0.0121 0.0017 0.5429 0.0123 -0.0027
MobileSAM 0.5622 0.4020 0.5126 0.4059 0.0000 0.5136 0.4067 0.0002 0.5196 0.4076 0.0004
MobileSAM+dropout 0.5486 0.4011 0.5220 0.4057 0.0002 0.5188 0.4065 0.0002 0.5263 0.4074 0.0004
MobileSAM+ensemble 0.5168 0.4012 0.5128 0.4210 0.0009 0.4802 0.4171 0.0013 0.4792 0.4070 0.0013
MobileSAM+ensemble+dropout 0.4974 0.4004 0.5315 0.4223 0.0006 0.5212 0.4193 0.0007 0.5112 0.4086 0.0012
@@
Experiment B. In Table 2, we present the performance metrics on the ID test set, along with an OOD split focused exclusively on hands. Our analysis reveals that the primary source of variability in IoU scores arises from the choice of evaluation set rather than other factors such as the application of dropout or variations in model architecture. Notably, model ensembling leads to improved IoU scores, particularly evident in the replaced green screen dataset.
@@
Experiment C. The findings of the skin tone diversity investigation suggest that the initial data collection efforts are encouraging, but additional data is required to make a more substantial contribution to this area of study. This is illustrated in Figure 4. Specifically, Figure 4(c) demonstrates that, in proportional terms, the HAGS dataset provides greater coverage of darker skin tones compared to WH and HRC. However, in terms of absolute counts, HRC and WH exhibit broader skin tone representation. These results indicate that rebalancing the skin tone distributions in WH and HRC could address some diversity limitations within these datasets. Nevertheless, the preferred approach remains the collection of more diverse datasets that are representative of real-world use cases, ensuring a broader and more inclusive distribution of skin tones.
(a) Monk Skin Tone (MST) Scale (Monk, 2019).
(b) Histogram in raw counts of MST representation in the real world images from the HAGS (Sharma et al., 2024), HRC (Sajedi et al., 2022), and WH (Shilkrot et al., 2019) datasets.
(c) Normalized histogram in raw counts of MST representation in the real world images from the HAGS (Sharma et al., 2024), HRC (Sajedi et al., 2022), and WH (Shilkrot et al., 2019) datasets.
Figure 4: This figure provides a summary of the findings from Experiment C on skin tone representation. Analyzing raw frequency counts, the HRC dataset (Sajedi et al., 2022) demonstrates higher absolute representation of darker skin tones. However, when normalized to proportional representation, the HAGS dataset (Sharma et al., 2024) complements WH (Shilkrot et al., 2019) and HRC by exhibiting an increased focus on darker skin tone demographics, addressing gaps in representation evident in the other datasets.
@@ 5 Discussion, Conclusions, and Future Work
There are several areas in which this work can be improved. One significant enhancement would be to increase the diversity and quantity of images and labels, particularly by increasing the training set size to be comparable to other industrial domain datasets. Although the test subjects were diverse, the size of the studies limited the range of diversity that could be included. Future work should further improve diversity with a larger participant pool that includes various age ranges, genders, and skin tones. We currently employ two static camera angles, which, while providing some diversity, resulted in hands being predominantly located in predictable, centered portions of the image, as indicated by the pixel occupancy heat maps (see Figure 5). This is a noted, but not significant, limitation since the rigid form of the glovebox limits operators to working in a known and relatively small reachable workspace. The study does not utilize depth sensors to acquire accompanying depth data, nor does it record robot trajectory information. Although both of which could be valuable for advancing other human-robot interaction (HRI) applications, their exclusion simplifies the experimental set-up while mirroring glovebox configurations in the real world. Another limitation is that the participants utilized a single brand of glove. Although gloves are typically similar in color they can vary slightly; and more importantly, they tend to yellow with age. Finally, the method employed for classifying skin tone diversity in Experiment 3 has inherent limitations. While this method may give a general sense of dataset diversity, reliance on basic image processing techniques restricts both the scalability and robustness of the approach. Unfortunately, the Monk Skin Tone (MST) scale (Monk, 2019), while valuable as a standard, lacks an open-source classifier, limiting our analysis.
@@
{acks}
This manuscript has been approved for unlimited release and has been assigned LA-UR-24-25500 (Rev. 1). This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This work was supported by the Laboratory Directed Research and Development program of LANL under project number 20210043DR, C1582/CW8217, and 20250048DR. LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001).
{sm}
5.1 HAGS Datasheet
This section introduced our Hand and Glove Segmentation Dataset (HAGS) datasheet (Gebru et al., 2021). The HAGS dataset was collected by the Nuclear and Applied Robotics Group and funded by Los Alamos National Laboratory.
Hosting, Long Term Preservation, and Maintenance: The dataset will be maintained by Selma Wanna (Email: slwanna@utexas.edu) and hosted on the Texas Data Repository: https://dataverse.tdl.org/.
@@ 5.1 HAGS Datasheet
The dataset consists of 191 videos, totaling approximately 9 hours of footage, along with 1728 annotated in-distribution and out-of-distribution frames, all originating from 10 diverse participants. Furthermore, the dataset emphasizes participants and procedures with various characteristics, such as differing skin tones, arm orientations, camera perspectives, and task execution methods. The two joint assembly tasks within this dataset include the assembly of a block tower and the disassembly of a tool box. In all, the diversity presented enables the dataset to cover a broad range of real-world scenarios.
@@
There are 16 videos per test subject, from which sampled frames and annotations were created over 2 experiments, 4 variables, 2 camera angles, and 10 subjects.
The dataset includes 10 diverse participants and is self-contained; it does not include external sources or assets.
@@
There are labels provided only for the images we have sampled. It is possible for other users to sample more images using the provided script: https://github.com/sanatnair/Glovebox_Segmentation_Dataset_Tools. Thereafter, software like LabelStudio (Tkachenko et al., 2020) can assist in creating the new images’ ground truths.
@@
The dataset is self-contained barring the instances with green screens with overlaid images. These images were sourced from Duck Duck Go images using a script: https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset/blob/main/data/convert_scripts/replace_green.py. There is no guarantee that the specific images will remain constant over time. There is not an archival version of this portion of the dataset that we can distribute.
@@
An ethical review process was conducted. An IRB exemption was granted for this study. This information is viewable online at https://dataverse.tdl.org/file.xhtml?fileId=599918&version=1.0.
@@
All participants received an informed consent form: https://dataverse.tdl.org/file.xhtml?fileId=599917&version=1.0 and were aware of our study and data collection.
@@
The software for labeling the data is available from LabelStudio (Tkachenko et al., 2020).
Frames are labeled as left and right hand with LabelStudio software (Tkachenko et al., 2020) and the MobileSAM (Zhang et al., 2023) backend. Models were trained on frames compressed to 256x256 px. Ground truth masks and original footage data is available, while raw unprocessed data is private.
@@
This repository: https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset/tree/main links to our public repository, which will contain a running list of associated papers or publications.
@@
5.2 Overview of Hand Segmentation Datasets
@@
GTEA (Fathi et al., 2011) 2011 C GoPro Cooking Kitchen
ADL (Pirsiavash and Ramanan, 2012) 2012 C GoPro Daily Life Home
EDSH (Li and Kitani, 2013) 2013 C - Daily Life Home / Outdoors
Interactive Museum (Baraldi et al., 2014) 2014 C - Gesture Museum
EgoHands (Bambach et al., 2015) 2015 C Google Glass Manipulation Social Setting
Maramotti (Baraldi et al., 2015) 2015 C - Gesture Museum
UNIGE Hands (Betancourt et al., 2015) 2015 C GoPro Hero3+ Daily Life Various Daily Locations
GUN-71 (Rogez et al., 2015) 2015 CD Creative Senz3D Daily Life Home
RGBD Egocentric Action (Wan and Aggarwal, 2015) 2015 CD Creative Senz3D Daily Life Home
Fingerwriting in mid-air (Chang et al., 2016) 2016 CD Creative Senz3D Writing Office
Ego-Finger (Huang et al., 2016) 2016 C - Gesture Various Daily Locations
ANS able-bodied (Likitlersuang and Zariffa, 2018) 2016 C Looxie 2 Daily Life Home
UT Grasp (Cai et al., 2017) 2016 C GoPro Hero2 Manipulation -
GestureAR (Mohatta et al., 2017) 2017 C Nexus 6 and Moto G3 Gesture Various Backgrounds
EgoGesture (Wu et al., 2017) 2017 C - Gesture -
Egocentric hand-action (Xu et al., 2016) 2017 D Softkinetic DS325 Gesture Office
BigHand2.2M (Yuan et al., 2017) 2017 D Intel RealSense SR300 Gesture/Pose -
Desktop Action (Cai et al., 2019) 2018 C GoPro Hero 2 Daily Life Office
Epic Kitchens (Damen et al., 2018) 2018 C GoPro Cooking Kitchen
FPHA (Garcia-Hernando et al., 2018) 2018 CD Intel RealSense SR300 Gesture / Pose / Daily life Home
EYTH (Khan and Borji, 2018) 2018 C - - -
HandOverFace (Khan and Borji, 2018) 2018 C - - -
EGTEA+ (Li et al., 2020) 2018 C SMI wearable eye-tracker Cooking Kitchen
THU-READ (Tang et al., 2019) 2018 CD Primesense Carmine Daily Life Home
EgoGesture (Zhang et al., 2018) 2018 CD Intel RealSense SR300 Gesture -
HOI (Kang et al., 2018) 2018 D Kinect V2 Manipulation -
EgoDaily (Cruz and Chan, 2019) 2019 C GoPro Hero5 Daily Life Various Daily Locations
ANS SCI (Likitlersuang et al., 2019) 2019 C GoPro Hero4 Daily Life Home
KBH (Wang et al., 2019) 2019 C HTC Vive Manipulation / Keyboard Office
WorkingHands (Shilkrot et al., 2019) 2019 CD Kinect V2 Assembly Industrial
Freihand (Zimmermann et al., 2019) 2019 C - Grasp / Pose Green Screen and Outdoors
MECCANO (Ragusa et al., 2020) 2020 C Intel RealSense SR300 Assembly Desk
HRC (Sajedi et al., 2022) 2022 C iPhone 11 Pro Assembly Industrial
HaDR (Grushko et al., 2023) 2023 CD Realsense L515 Manipulation Industrial